OpenStudy (jamesreed30):

Find the exact circumference in terms of pi of a circle with the given radius. 5 inches

OpenStudy (ovidia):

given: r = 5 C = 2(pi)r = 2 *3.14*5 = 31.4 in

OpenStudy (jamesreed30):

That's what I got also, but it says exact answer and what does that mean?

OpenStudy (ovidia):

C = 2πr = πd --- (where r = radius, d = diameter = 2 times the radius) C = r * (2π) = r * (6.283) --- (π = Greek letter pi = 3.1415...) So, for your problem where the radius, r, is given as 5 inches: C = 5 * (2π) = 10π = 31.415... inches --- (since this number goes on indefinitely, "10π inches" is about the most exact you can get without rounding)

OpenStudy (mrnood):

it says "exact answer in terms of pi" this means oyu do NOT evaluate pi the circumference is 2 pi r in this case C= 10π